# TensorHyper-VQC: A Tensor-Train-Guided Hypernetwork for Robust and Scalable Variational Quantum Computing

### v2

@@ 2.5 Application to Combinatorial Optimization: Max-Cut
We next evaluate the proposed TensorHyper-VQC architecture on a paradigmatic combinatorial optimization problem, which is the Max-Cut problem, formulated within the classical quantum approximate optimization algorithm (QAOA) framework [wang2018quantum]. The Max-Cut problem seeks a bipartition of a graph that maximizes the number of edges connecting the two partitions. Formally, for a graph $G=(V,E)$ with vertex set $V=V_{1}\cup V_{2}$, the objective is to maximize the number (or total weight) of edges between $V_{1}$ and $V_{2}$.
@@
In our experiments, TensorHyper-VQC employs a compact TT network to generate all variational circuit parameters required by QAOA across different graph instances. The TT network is constructed with an input dimension of $10$, output dimensions $[1,1]$, and TT-ranks $[1,4,1]$, resulting in a lightweight parameterization with only $41$ trainable parameters (including bias). We benchmarked the approach on 10 randomly generated 20-qubit Erdos-Renyi graphs [engel2004large], where each edge is included independently with probability $0.5$.
@@
Comparison Against Noise-Mitigation-Enhanced QAOA. To assess performance in realistic conditions, we further compare TensorHyper-VQC with classical QAOA enhanced with ZNE and REM, two widely used quantum error mitigation techniques. The results in Table 3 demonstrate that TensorHyper-VQC still outperforms QAOA (ZNE+REM), with an average improvement of +4.05+4.05 (or $16.09\%$). This shows that TensorHyper-VQC achieves competitive or better results without relying on quantum noise mitigation techniques, underscoring its inherent noise resilience and robust parameterization.
@@
Table 4: Experimental results on a $156$-qubit IBM quantum processor (Heron r2) comparing Classical QAOA with ZNE+REM noise mitigation and TensorHyper-VQC on the MaxCut problem over 10 random graph instances ($20$-Qubit Graphs).
@@ 2.6 Application to Quantum Simulation: Molecular Hamiltonians
To further demonstrate the practical utility of the proposed TensorHyper-VQC framework, we evaluated its performance on a canonical quantum simulation task: the ground-state energy estimation of the lithium hydride (LiH) molecule. Quantum chemistry simulation is widely regarded as one of the most compelling near-term applications of variational quantum algorithms. It thus serves as a stringent benchmark for assessing the scalability, expressivity, and robustness of variational architectures. In this study, we considered a 4-qubit reduced Hamiltonian of LiH, obtained by freezing core orbitals and truncating high-lying virtual orbitals on an STO-3G basis. This reduced model preserves the essential many-body correlations while enabling tractable simulations within the capabilities of current quantum hardware.
We configured TensorHyper-VQC using a TT hypernetwork to generate the variational ansatz parameters. The ansatz consists of $L=2$ layers of parameterized single-qubit rotations ($R_{X},R_{Y},R_{Z}$) followed by a ring of CNOT entanglers, resulting in a 24-parameter variational circuit in the conventional VQE setting. By contrast, the TT hypernetwork compactly represents this parameter vector using input dimensions $[4,6]$ and TT-ransk $[1,2,1]$, thereby reducing the number of trainable parameters to only $9$ without compromising expressivity. The TT parameters were optimized classically using the COBYLA algorithm [regis2011stochastic] to minimize the ground-state energy of the LiH Hamiltonian.
@@
Validation on Real Quantum Hardware. To assess the framework’s robustness under practical hardware constraints, we further implemented the LiH simulation on a $156$-qubit IBM Quantum processor (Heron r2). As shown in Table 6, TensorHyper-VQC once again outperformed the baseline. The experimentally measured ground-state energy was 7.835377-7.835377 (+0.026752+0.026752 Ha error), compared with 7.809882-7.809882 Ha (+0.052247+0.052247 Ha error) achieved by the conventional VQE with ZNE+REM. The significant reduction in error on real hardware underscores the advantage of TT-guided parameterization in mitigating the effects of noise, decoherence, and device imperfections.
@@
Overall, these findings demonstrate that TensorHyper-VQC attains near-chemical-accuracy ground-state energy estimates with dramatically fewer parameters and simultaneously exhibits intrinsic robustness to quantum noise and hardware imperfections. The structured low-rank parameterization produces ansätze that are both expressive and less sensitive to perturbations, resulting in superior performance in both simulated and experimental quantum-simulation settings.
@@ 3 Discussion
In this work, we introduced TensorHyper-VQC, a novel tensor-train-guided hypernetwork architecture specifically designed to overcome two fundamental challenges in VQC: barren plateaus and susceptibility to quantum noise. By strategically decoupling parameter optimization from quantum hardware and delegating it to a classical TT network, TensorHyper-VQC substantially enhances the scalability, robustness, and practicality of variational quantum algorithms. This classical-quantum decoupling reshapes the optimization landscape, improving gradient flow and mitigating the detrimental effects of noise on parameter updates.
Conventional VQCs often suffer from vanishing gradients and noise-induced instability, limiting their expressivity and applicability to realistic quantum devices. Numerous recent studies have attempted to address these challenges through advanced parameterization strategies [larocca2023theory] and structural optimizations [holmes2022connecting]. Low-rank tensor decompositions, in particular, have emerged as a practical approach to reduce parameter dimensionality and enhance generalization in both classical and quantum settings. Building on these advances, TensorHyper-VQC leverages an entirely classical TT hypernetwork to generate variational parameters, thereby combining the representational power of tensor networks with the trainability benefits of low-rank optimization.
This architectural shift yields several advantages. First, by transferring all optimization tasks to the classical domain, TensorHyper-VQC circumvents the stochasticity introduced by quantum measurement, leading to more stable and efficient training. Second, the inherent low-rank structure of the TT parameterization introduces beneficial inductive biases, such as noise averaging and variance reduction across TT cores, which directly contribute to noise resilience. Our theoretical analyses, based on NTK theory and variance reduction, further substantiate these benefits by demonstrating that the variance of quantum noise on TT-core gradients decreases with increasing qubit count —an effect not exploited by conventional VQCs.
@@ 4.1 An Introduction to Neural Tangent Kernel Theory
The Neural Tangent Kernel (NTK) [bietti2019inductive, jacot2018neural, liu2022representation] is a potent theoretical framework that illustrates the efficient trainability of deep neural networks despite their inherent complexity. Fundamentally, the NTK theory examines how marginal changes in a network’s parameters affect its predictions. A network is deemed ”well-conditioned,” leading to smoother training and faster convergence, when minor parameter adjustments reliably yield stable, predictable shifts in output.
@@
Our theoretical framework for TensorHyper-VQC directly leverages NTK theory to analyze its trainability. The VQC and the TT network determine the NTK matrix for TensorHyper-VQC. A key theoretical finding is that the smallest eigenvalue of TensorHyper-VQC’s NTK is strictly larger than that of a conventional VQC. This indicates a better-optimized landscape for TensorHyper-VQC, leading to faster, more stable convergence by mitigating the vanishing gradient problem. Thus, the structured parameterization afforded by the TT network directly contributes to improved trainability as supported by our NTK analysis.
@@ 4.2 Tensor-Train Network and TTN+VQC
The TT network [oseledets2011tensor], also known as the matrix product state [cirac2021matrix], is one of the most widely used tensor network architectures in both machine learning and quantum many-body physics. It provides an efficient way to represent and manipulate high-dimensional data by decomposing a large tensor into a sequence of lower-order tensors (called TT cores) arranged in a one-dimensional chain.
@@
When integrated with VQC, the TT network forms a hybrid quantum-classical architecture, traditionally referred to as TTN+VQC [qi2023theoretical]. In this framework, the TT network typically serves as a classical pre-processing or feature-compression module that reduces the practical input dimension before quantum encoding. The VQC then processes the compressed features, leveraging quantum superposition and entanglement to enhance the model’s expressivity and capture nonlinear correlations that are difficult to represent classically. This complementary interplay positions TTN+VQC as a promising paradigm for tackling complex and high-dimensional quantum machine learning tasks.
@@ 4.3 Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm (QAOA) [zhou2020quantum, zhu2022adaptive] is a hybrid quantum-classical algorithm designed to find approximate solutions to combinatorial optimization problems. It iteratively optimizes the parameters of a quantum circuit to maximize an objective function defined by the problem. One common application of QAOA is solving the Max-Cut problem. The Max-Cut problem seeks to partition the vertices of a graph into two sets, maximizing the number of edges connecting vertices in different sets.
@@ 4.4 Variational Quantum Eigensolver
The Variational Quantum Eigensolver (VQE) [kandala2017hardware, zhang2022variational] is a prominent hybrid quantum-classical algorithm designed to find the ground state energy of a given Hamiltonian. It works by variably optimizing the parameters of a quantum circuit to prepare a quantum state. The expectation value of the Hamiltonian is then measured for this state on a quantum computer and fed back to a classical optimizer. The classical optimizer then updates the quantum circuit’s parameters to minimize energy, iteratively searching for the quantum state corresponding to the lowest possible energy, i.e., the ground state. VQE is a promising algorithm for NISQ devices due to its relatively shallow circuit depth requirements.
@@ 4.5 TT Network Input and Feature Processing
In the TensorHyper-VQC framework, the classical TT network is a hypernetwork that maps problem-specific classical features to the full quantum circuit parameters required for the target VQC task. The TT network receives a vectorized representation of the charge stability diagram as input for classifying quantum dots. Specifically, each 50 × 50 image is flattened into a 2,500-dimensional real-valued vector, which is then projected or reshaped as needed to match the input dimension structure of the TT-cores. For combinatorial optimization tasks, such as Max-Cut, the TT network input consists of concise task-level descriptors, including the degree histogram and other graph statistics.
@@ 4.8 The proof of Upper Bound on Approximation Capability
As for the first term $\lVert h^{*}(\textbf{x})-f_{\textbf{w}^{*}}(\textbf{x})\rVert_{1}$, it is related to the expresssivity limitation of VQC. Based on our derived theoretical results in [qi2023theoretical], for an $L$-depth and $U$-qubit VQC, we have:
@@ 4.10 Proof of Robustness to Quantum Noise
The result demonstrates that both the bias and variance shrink with system size: the bias decreases as $\frac{1}{\sqrt{UL}}$ and the variance as $\frac{1}{UL}$. Even systematic measurement errors become less harmful as the circuit grows.
@@
which means that the noise variance and bias are independent of $U$ or $L$. As circuits grow, noise accumulates instead of averaging out. By contrast, in TensorHyper-VQC, the Jacobian naturally filters the noise. This gives the system an inherent robustness that improves with scale.
@@ 6 Data Availability Statement
The dataset used in our experiments on quantum dot classification can be downloaded from the website: https://gitlab.com/QMAI/mlqe_\_2023_\_edx.

### v3

@@ 2.5 Application to Combinatorial Optimization: Max-Cut
We next evaluate the proposed TensorHyper-VQC architecture on a paradigmatic combinatorial optimization problem, which is the Max-Cut problem, formulated within the classical quantum approximate optimization algorithm (QAOA) framework [45]. The Max-Cut problem seeks a bipartition of a graph that maximizes the number of edges connecting the two partitions. Formally, for a graph $G=(V,E)$ with vertex set $V=V_{1}\cup V_{2}$, the objective is to maximize the number (or total weight) of edges between $V_{1}$ and $V_{2}$.
@@
In our experiments, TensorHyper-VQC employs a compact TT network to generate all variational circuit parameters required by QAOA across different graph instances. The TT network is constructed with an input dimension of $10$, output dimensions $[1,1]$, and TT-ranks $[1,4,1]$, resulting in a lightweight parameterization with only $41$ trainable parameters (including bias). We benchmarked the approach on 10 randomly generated 20-qubit Erdos-Renyi graphs [12], where each edge is included independently with probability $0.5$.
@@
Comparison Against Noise-Mitigation-Enhanced QAOA. To assess performance under realistic conditions, we further compare TensorHyper-VQC with classical QAOA enhanced by ZNE and REM, two widely used quantum error-mitigation techniques. The results in Table 3 demonstrate that TensorHyper-VQC still outperforms QAOA (ZNE+REM), with an average improvement of +4.05+4.05 (or $16.09\%$). This shows that TensorHyper-VQC achieves competitive or better results without relying on quantum noise mitigation techniques, underscoring its inherent noise resilience and robust parameterization.
@@
Table 4: Experimental results on a $156$-qubit IBM quantum processor (Heron r2) comparing Classical QAOA with ZNE+REM noise mitigation and TensorHyper-VQC on the MaxCut problem over 10 random graph instances (2020-Qubit Graphs).
@@ 2.6 Application to Quantum Simulation: Molecular Hamiltonians
To further demonstrate the practical utility of the proposed TensorHyper-VQC framework, we evaluated its performance on a canonical quantum simulation task: the ground-state energy estimation of the lithium hydride (LiH) molecule. Quantum chemistry simulation is widely regarded as one of the most compelling near-term applications of variational quantum algorithms. It thus serves as a stringent benchmark for assessing the scalability, expressivity, and robustness of variational architectures. In this study, we considered a 4-qubit reduced Hamiltonian of LiH, obtained by freezing core orbitals and truncating high-lying virtual orbitals on an STO-3G basis. This reduced model preserves the essential many-body correlations while enabling tractable simulations on current quantum hardware.
We configured TensorHyper-VQC using a TT hypernetwork to generate the variational ansatz parameters. The ansatz consists of $L=2$ layers of parameterized single-qubit rotations ($R_{X},R_{Y},R_{Z}$) followed by a ring of CNOT entanglers, resulting in a 24-parameter variational circuit in the conventional VQE setting. By contrast, the TT hypernetwork compactly represents this parameter vector using input dimensions $[4,6]$ and TT-ransk $[1,2,1]$, thereby reducing the number of trainable parameters to only $9$ without compromising expressivity. The TT parameters were optimized classically using the COBYLA algorithm [37] to minimize the ground-state energy of the LiH Hamiltonian.
@@
Validation on Real Quantum Hardware. To assess the framework’s robustness under practical hardware constraints, we further implemented the LiH simulation on an IBM Quantum $156$-qubit processor (Heron r2). As shown in Table 6, TensorHyper-VQC once again outperformed the baseline. The experimentally measured ground-state energy was 7.835377-7.835377 (+0.026752+0.026752 Ha error), compared with 7.809882-7.809882 Ha (+0.052247+0.052247 Ha error) achieved by the conventional VQE with ZNE+REM. The significant reduction in error on real hardware underscores the advantage of TT-guided parameterization in mitigating the effects of noise, decoherence, and device imperfections.
@@
Overall, these findings demonstrate that TensorHyper-VQC attains near-chemical-accuracy ground-state energy estimates with dramatically fewer parameters and simultaneously exhibits intrinsic robustness to quantum noise and hardware imperfections. The structured low-rank parameterization yields ansätze that are both expressive and less sensitive to perturbations, leading to superior performance in both simulated and experimental quantum-simulation settings.
@@ 3 Discussion
In this work, we introduce TensorHyper-VQC, a novel tensor-train-guided hypernetwork architecture designed to address two fundamental challenges in VQC: barren plateaus and susceptibility to quantum noise. By strategically decoupling parameter optimization from quantum hardware and delegating it to a classical TT network, TensorHyper-VQC substantially enhances the scalability, robustness, and practicality of variational quantum algorithms. This classical-quantum decoupling reshapes the optimization landscape, improving gradient flow and mitigating the detrimental effects of noise on parameter updates.
Conventional VQCs often suffer from vanishing gradients and noise-induced instability, limiting their expressivity and applicability to realistic quantum devices. Numerous recent studies have attempted to address these challenges through advanced parameterization strategies [23] and structural optimizations [16]. Low-rank tensor decompositions, in particular, have emerged as a practical approach to reduce parameter dimensionality and enhance generalization in both classical and quantum settings. Building on these advances, TensorHyper-VQC leverages an entirely classical TT hypernetwork to generate variational parameters, thereby combining the representational power of tensor networks with the trainability benefits of low-rank optimization.
This architectural shift yields several advantages. First, by transferring all optimization tasks to the classical domain, TensorHyper-VQC mitigates the stochasticity introduced by quantum measurement, yielding more stable and efficient training. Second, the inherent low-rank structure of the TT parameterization introduces beneficial inductive biases, such as noise averaging and variance reduction across TT cores, which directly contribute to noise resilience. Our theoretical analyses, based on NTK theory and variance reduction, further substantiate these benefits by demonstrating that the variance of quantum noise in TT-core gradients decreases with increasing qubit count—an effect not exploited by conventional VQCs.
@@ 4.1 An Introduction to Neural Tangent Kernel Theory
The Neural Tangent Kernel (NTK) [3, 19, 25] is a potent theoretical framework that illustrates the efficient trainability of deep neural networks despite their inherent complexity. Fundamentally, the NTK theory examines how marginal changes in a network’s parameters affect its predictions. A network is deemed ”well-conditioned,” leading to smoother training and faster convergence, when minor parameter adjustments reliably yield stable, predictable output shifts.
@@
Our theoretical framework for TensorHyper-VQC directly leverages NTK theory to analyze its trainability. The VQC and the TT network determine the NTK matrix for TensorHyper-VQC. A key theoretical finding is that the smallest eigenvalue of the NTK of TensorHyper-VQC is strictly greater than that of a conventional VQC. This indicates a more optimized landscape for TensorHyper-VQC, leading to faster, more stable convergence by mitigating the vanishing-gradient problem. Thus, the structured parameterization afforded by the TT network directly improves trainability, as supported by our NTK analysis.
@@ 4.2 Tensor-Train Network and TTN+VQC
The TT network [32], also known as the matrix product state [9], is one of the most widely used tensor network architectures in both machine learning and quantum many-body physics. It provides an efficient way to represent and manipulate high-dimensional data by decomposing a large tensor into a sequence of lower-order tensors (called TT cores) arranged in a one-dimensional chain.
@@
When integrated with VQC, the TT network forms a hybrid quantum-classical architecture, traditionally referred to as TTN+VQC [35]. In this framework, the TT network typically serves as a classical pre-processing or feature-compression module that reduces the practical input dimension before quantum encoding. The VQC then processes the compressed features, leveraging quantum superposition and entanglement to enhance the model’s expressivity and capture nonlinear correlations that are difficult to represent classically. This complementary interplay positions TTN+VQC as a promising paradigm for tackling complex and high-dimensional quantum machine learning tasks.
@@ 4.3 Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm (QAOA) [49, 50] is a hybrid quantum-classical algorithm designed to find approximate solutions to combinatorial optimization problems. It iteratively optimizes the parameters of a quantum circuit to maximize an objective function defined by the problem. One common application of QAOA is solving the Max-Cut problem. The Max-Cut problem seeks to partition the vertices of a graph into two sets, thereby maximizing the number of edges between the two sets.
@@ 4.4 Variational Quantum Eigensolver
The Variational Quantum Eigensolver (VQE) [21, 48] is a prominent hybrid quantum-classical algorithm designed to find the ground state energy of a given Hamiltonian. It works by variably optimizing the parameters of a quantum circuit to prepare a quantum state. The expectation value of the Hamiltonian is then measured for this state on a quantum computer and fed back to a classical optimizer. The classical optimizer then updates the quantum circuit’s parameters to minimize energy, iteratively searching for the quantum state corresponding to the lowest possible energy, i.e., the ground state. VQE is a promising algorithm for NISQ devices due to its relatively shallow circuit depth requirements.
@@ 4.5 TT Network Input and Feature Processing
In the TensorHyper-VQC framework, the classical TT network is a hypernetwork that maps problem-specific classical features to the full quantum circuit parameters required for the target VQC task. The TT network takes as input a vectorized representation of the charge stability diagram to classify quantum dots. Specifically, each 50 × 50 image is flattened into a 2,500-dimensional real-valued vector, which is then projected or reshaped as needed to match the input dimension structure of the TT-cores. For combinatorial optimization tasks, such as Max-Cut, the TT network input consists of concise task-level descriptors, including the degree histogram and other graph statistics.
@@ 4.8 The proof of Upper Bound on Approximation Capability
As for the first term $\lVert h^{*}(\textbf{x})-f_{\textbf{w}^{*}}(\textbf{x})\rVert_{1}$, it is related to the expresssivity limitation of VQC. Based on our derived theoretical results in [35], for an $L$-depth and $U$-qubit VQC, we have:
@@ 4.10 Proof of Robustness to Quantum Noise
The result demonstrates that both bias and variance shrink with system size: the bias decreases as $\frac{1}{\sqrt{UL}}$, and the variance decreases as $\frac{1}{UL}$. Even systematic measurement errors become less harmful as the circuit grows.
@@
which means that the noise variance and bias are independent of $U$ or $L$. As circuits scale, noise accumulates rather than averaging out. By contrast, in TensorHyper-VQC, the Jacobian naturally filters the noise. This gives the system an inherent robustness that improves with scale.
@@ 6 Data Availability Statement
The dataset used in our experiments on quantum dot classification is available at https://gitlab.com/QMAI/mlqe_\_2023_\_edx.